# HyperPg - Prototypical Gaussians on the Hypersphere for Interpretable Deep Learning

### v2

@@ 3.2 Gaussian Prototypes
This formulation as a PDF has several advantages: A) The similarity can be interpreted as the likelihood of being sampled from the Gaussian prototypes, which is more meaningful than a distance metric in a high-dimensional latent space. B) Prototypes can adapt their shape using a full covariance matrix, allowing different variances along various feature dimensions, offering more flexibility in shaping the latent space. However, this increases computational requirements, especially with EM clustering. Gaussians prototypes lose the advantages of hyperspherical prototypes for classification and regression Mettes et al. (2019).
3.3 Gaussian Prototypes on the Hypersphere - HyperPg
Prototypical Gaussians on the Hypersphere (HyperPg) combine the advantages of Gaussian and hyperspherical prototypes. HyperPg prototypes are defined as ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ with a directional anchor vector $\bm{{\alpha}}$, scalar mean similarity $\mu$ and scalar standard deviation (std) $\sigma$. HyperPg prototypes learn a $1D$ Gaussian distribution over the cosine similarities to the anchor vector $\bm{{\alpha}}$. Because the cosine similarity is bounded to $[-1,1]$, HyperPg’s similarity measure is defined as the PDF of the truncated Gaussian distribution within these bounds. Let $\mathcal{G}(x,\mu,\sigma)$ be the cumulative Gaussian distribution function. Then, HyperPg’s similarity measure based on the truncated Gaussian distribution is defined as
@@ 3.3 Gaussian Prototypes on the Hypersphere - HyperPg
Figure 2: HyperPg learns a Gaussian distribution with mean $\mu$ and std $\sigma$ of cosine similarities with the anchor $\bm{{\alpha}}$. The plots visualize HyperPg’s activations on the $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, std $\sigma=0.1$ and and various mean cosine similarities $\mu\in[-1,1]$ for 10k random samples. The anchor $\bm{{\alpha}}$ is shown as a red arrow, and std $\sigma$ governs the width of the distribution. For $\mu=1$, the distribution is aligned with the vector; for $\mu=-1$ it is on the opposite side of the hypersphere. Due to the cosine similarity, setting $11>\mu>-$ 1 interpolates the area of highest probability density between both poles, leading to a ring shape on the hypersphere’s surface.
@@
Fig. 2 illustrates the activations of HyperPg’s similarity function on the surface of a $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, fixed std $\sigma=0.1$ and various mean values $\mu\in[-1,1]$. The anchor $\bm{{\alpha}}$ defines a prototypical direction vector in latent space $\mathcal{Z}$, similar to other hyperspherical prototypes, and is visualized as a red arrow. The learned Gaussian distribution of cosine similarities is projected onto the hypersphere’s surface with std $\sigma$ governing the spread of the distribution, and mean $\mu$ the expected distance to the anchor $\bm{{\alpha}}$. For $\mu=1$, the distribution centers around the anchor, as the cosine similarity is $1$ if two vectors point in the same direction. For $\mu=-1$, the distribution is the on opposite side of the hypersphere, as the cosine similarity is $-1$ for vectors pointing in opposite directions.
@@
For values of $11>\mu>-$ 1, the distribution forms a hollow ring around the anchor vector $\bm{{\alpha}}$. This occurs because the cosine similarity for these $\mu$ values expects the activating vectors to point in a different direction than the anchor, without specifying the direction. Imagining the interpolation between $\mu=1$ and $\mu=-1$, the probability mass moves from one pole of the hypersphere to the other, stretching like a rubber band over the surface. For $\mu=0$, the expected cosine similarity indicates that vectors with the highest activation are orthogonal to the anchor $\bm{{\alpha}}$. Since no specific direction is indicated, the entire hyperplanar segment orthogonal to the anchor has the highest activation. This activation pattern for the cosine similarity would typically require an infinite mixture of prototype vectors pointing in all directions in this hyperplane. HyperPg achieves the same effect by learning only one prototype vector (the anchor) and just two additional scalar parameters. This significantly increases HyperPg’s representational power compared to standard hyperspherical and Gaussian prototypes and is a major difference to the von Mises-Fisher distribution (see also Appendix A).
@@
Combining the strengths of hyperspherical and Gaussian prototypes, HyperPg can learn more complex structures in the latent space, such as human-defined concepts, and provide a more meaningful similarity measure. HyperPg can be easily adapted to other probability distributions with additional desirable properties. Possible candidate distributions are elaborated on in Appendix A. Similarly, it is possible to exchange the cosine similarity to other similarity measures or functions, and learn an untruncated PDF over their output, making the HyperPg idea transferable to applications outside of prototype learning.
@@ 4 Training
The original ProtoPNet implementation uses three loss functions: a task specific loss like crossentropy for classification, a cluster loss to increase compactness within a class’s cluster, and a separation loss to increase distances between different prototype clusters. However, the prototypes of ProtoPNet and its successors are learned in a black-box manner.
HyperPgNet is a new inherently interpretable deep learning approach built on HyperPg prototypes. It introduces a “Right for the Right Concept” loss, inspired by “Right for the Right Reasons” (Ross et al., 2017), to restrict the learned prototypes to human-defined concepts. This focus enhances the interpretability and minimizes the influence of confounding factors. This section first provides an overview of the used prototype learning losses, then introduces the Right for the Right Concept loss, and finally discusses the overall network architecture and final multi-objective loss.
@@ 4.1 Prototype Losses
where $s(\cdot,\cdot)$ is some similarity measure. The $L_{\mathrm{Clst}}$-Loss function increases compactness by increasing the similarity between prototypes ${\bm{p}}_{c}$ of class $c$ latent embeddings ${\bm{z}}_{c}$ of class $c$ over all samples.
@@
The $L_{\mathrm{Sep}}$ function punishes high similarity values between a latent vector ${\bm{z}}_{c}$ of class c and prototypes ${\bm{p}}_{\neg c}$ not belonging to $c$, thereby separating the clusters in latent space. Please note, ProtoPNet (Chen et al., 2019) use a slightly different notation by working with the $L_{2}$ distance, instead of a similarity measure.
In HyperPgNet the learned prototypes are not class exclusive, but shared among different classes. Instead, each prototype is assigned to one human-defined concept $k\in K$. The Concept Activation Regions (CAR, Crabbé & van der Schaar, 2022) model proposes a concept density loss defined as
$L_{Density}$ $=-\frac{1}{|K|}\sum_{k\in K}\frac{1}{|{\bm{P}}_{k}|}\sum_{{\bm{p}}_{k}\in{\bm{P}}_{k}}\phi({\bm{p}}_{k},{\bm{Z}}_{k})-\phi({\bm{p}}_{k},{\bm{Z}}_{\neg k})$ (8)
$\phi({\bm{p}},{\bm{Z}})$ $=\frac{1}{|{\bm{Z}}|}\sum_{{\bm{z}}\in{\bm{Z}}}s({\bm{p}},{\bm{z}}).$ (9)
The density loss uses the similarity aggregation $\phi({\bm{p}},{\bm{Z}})$ which computes the mean response of a prototype ${\bm{p}}$ with a set of latent features ${\bm{Z}}$. The density loss therefore computes the mean response of correctly assigned prototypes minus the mean response of incorrectly assigned prototypes. The loss functions proposed by ProtoPNet compute the maximum response of correctly and incorrectly assigned prototypes instead.
4.2 Right Concept Loss
To ensure prototypes actually correspond to the input pixels containing the concept, and do not respond to other factors in the background, HyperPgNet introduces the “Right for the Right Concept” (RRC) -Loss inspired by “Right for the Right Reasons” (RRR, Ross et al., 2017). The RRC loss is defined as
$L_{\mathrm{RRC}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{k\in K}\left({\bm{A}}_{{\bm{x}}_{i},k}\frac{\partial}{\partial{\bm{x}}_{i}}\sum_{{\bm{p}}_{k}\in{\bm{P}}_{k}}s\left({\bm{p}}_{k},\mathrm{Enc}\left({\bm{x}}_{i}\right)\right)\right)^{2},$ (10)
with binary annotation matrix ${\bm{A}}_{{\bm{x}}_{i},k}\in\{0,1\}^{N\times\times W\times H}$ for each input sample ${\bm{x}}_{i}$ and concept $k$. This annotation matrix defines for each input image, which pixels contain which concept. In the original RRR paper, the annotation matrix specified relevant regions for the classification task, steering the model’s activations away from confounding factors in the background. In HyperPgNet the RRC-Loss further strengthens the prototype-concept association.
@@ 4.3 Multi-Objective Loss Function
$L=L_{\mathrm{CE}}+\lambda_{\mathrm{Clst}}L_{\mathrm{Clst}}+\lambda_{\mathrm{Sep}}L_{\mathrm{Sep}}+\lambda_{\mathrm{RRC}}L_{\mathrm{RRC}},$
where $L_{\mathrm{CE}}$ is the cross-entropy loss over network predictions and ground truth image labels. ProtoPNet uses $\lambda_{\mathrm{Clst}}=0.8$ and $\lambda_{\mathrm{Sep}}=0.08$ (Chen et al., 2019). For HyperPgNet, the different loss terms are weighted equally, meaning all $\lambda=1$.
@@ 4.4 Network Architecture
Figure 3: HyperPgNet Architecture. The HyperPg module can be easily exchanged to other prototype formulations such as ProtoPNet. HyperPgNet uses the truncated Gaussian distribution as density estimator, but other PDFs are possible.
Fig. 3 illustrates HyperPgNet’s Architecture for interpretable image classification. HyperPgNet uses a pretrained feature encoder such as ConvNext-tiny (Liu et al., 2022) or MiT-B4 (Xie et al., 2021) as a backbone model. A neck consisting of two $1\times 1$ convolution layers with ReLU activation in between projects the high dimensional feature map of the backbone into a lower dimensional feature space.
The HyperPg prototype module is implemented with two layers: First, a prototype learning layer computes the cosine similarity of the learnable HyperPg anchors $\bm{{\alpha}}$ to the latent vectors produced by the neck. Second, a Density Estimation layer with learnable parameters mean $\mu$ and std $\sigma$ computes the Gaussian PDF over the activations of the previous layer. Future extensions can adapt both components independently, for example by implementing a hyperbolic similarity measure or a multi-modal probability distribution.
Finally, the output of the HyperPg Module is passed through a single fully connected layer (FCL) to produce the output logits or class scores.
5 Concept Extraction at Scale
To learn prototypes based on human-defined concepts with RRC-Loss for image classification, pixel level annotations are required. Foundation models are leveraged to generate these annotations, reducing the need for human labeling.
The CUB-200-2011 (Wah et al., 2011) annotates each image with up to 15 part locations (e.g., head, tail, wing). Each part is labeled with some attribute describing the coloring, pattern or shape. A Segment Anything 2 (SAM2) model (Ravi et al., 2024) creates pixel-level segmentation masks based on the part locations.
The Stanford Cars dataset (Krause et al., 2013) does not provide part level annotations. We defined a list of 10 car parts (e.g., wheel, headlight, radiator) and use a Grounding DINO model (Liu et al., 2023) to generate bounding boxes for each part. Finally, SAM2 creates pixel-level segmentation masks based on the bounding boxes. This automated pipeline labeled the 16k images of Stanford Cars within 2h on a NVIDIA 4060Ti with 16GB VRAM, and can be easily adapted to other datasets. Fig. 4 presents some examples for CUB-200-2011 and Stanford Cars.
(a) Concept Segmentation Masks for CUB-200-2011 based on provided positive and negative part annotations.
(b) Concept Segmentation Mask for Stanford Cars generated from a list of human-defined car parts.
Figure 4: Examples for Automated Concept Extraction.
@@ 6 Experiments
The experiments were performed on two datasets: CUB-200-2011 (CUB) with 200 bird species Wah et al. (2011) and Stanford Cars (CARS) with 196 car models Krause et al. (2013). Implementation details such as data preprocessing or model implementation are provided in Appendix B.
Birds Cars
# Prototypes MiT-B4 ConvNeXt # Prototypes MiT-B4 ConvNeXt
BB Baseline - 17.7 74.2 - 1.9 57.5
CBM - 75.7 77.9 - 79.9 81.0
PP ProtoPNet (PPN) 2000 68.0 68.1 1960 86.4 87.0
PPN + HyperPG 2000 70.5 65.0 1960 87.4 77.3
CAP HyperPgNet $-L_{\mathrm{RRC}}$ 300 76.5 76.9 180 88.6 88.9
(Ours) HyperPgNet $+L_{\mathrm{RRC}}$ 300 74.1 71.5 180 81.2 86.0
Table 1: Test Top-1 Accuracy. Models are grouped by their reasoning process. BB: Black Box. PP: Prototypical Parts. CAP: Concept Aligned Prototypes.
@@ 6.1 Quantitative Results
Figure 5: CUB Test Accuracy per Epoch.
Table 1 reports the top-1 accuracy for the tested models on the CUB and CARS datasets with different backbones. HyperPgNet achieves comparable results with the black box CBM and baseline models. Notably, HyperPgNet outperforms CBM on CARS, as it can differentiate between different modes for the same concept by learning multiple prototypes per concept. CBM is restricted to a binary concept presence prediction. Including ${\mathcal{L}}_{\mathrm{RRC}}$ slightly lowers HyperPgNet’s performance but is still comparable to the black-box models.
Fig. 5 shows the test accuracy over epochs for CUB with ConvNeXt backbone. The concept aligned models CBM and HyperPgNet are the fastest to converge at around 100 epochs. ProtoPNet requires over 400 training epochs until the best performance level is reached. However, switching from the standard $L_{2}$ based prototypes to HyperPg prototypes effectively halves the required training epochs.
6.2 Qualitative Comparison
Fig. 7 shows the pixel attribution of the highest activated prototype on each image for the tested prototype learning models with ConvNeXt backbone. In contrast to prior work (e.g., Chen et al., 2019; Ukai et al., 2023), this visualization is based on the prototype gradients, similar to saliency maps such as GradCAM (Selvaraju et al., 2017), which should avoid spatial misalignment often observed in prototype explanations (Sacha et al., 2024). The concept aligned prototypes learned by HyperPgNet focus more on the relevant parts of the image subject than the class-based prototypes of ProtoPNet. The network seems to focus correctly on either the overall color of the bird and ignoring the background (first three examples) or on specific markings (last three examples). The addition of the ${\mathcal{L}}_{\mathrm{RRC}}$ loss seems to further increase the main focus of the prototype activation on the relevant image regions. For ProtoPNet, no qualitative difference can be detected between the standard $L_{2}$ or HyperPg prototypes. With both prototype formulations, ProtoPNet is likely to also activate on parts of the image background. This is a fundamental weakness of training ProtoPNet in an opaque manner without alignment to relevant concepts, which HyperPgNet addresses.
On the CARS dataset, the effect of ${\mathcal{L}}_{\mathrm{RRC}}$ is more apparent. The defined concepts for this dataset are limited to car parts without fine-grained differentiation between patterns or shapes, such as the concepts on CUB. For this dataset, one of the most predictive image regions across models seems to be the front bumper of the cars, which has not been annotated with its own concept. As shown in Fig. 6, the concept-aligned prototypes still activate in this region. However, by training the model with the ${\mathcal{L}}_{\mathrm{RRC}}$ loss, the prototypes focus more on the annotated image regions.
Figure 6: Concept prototype activation after training with and without ${\mathcal{L}}_{\mathrm{RRC}}$ with MiT-B4 backbone.
Figure 7: Highest prototype activation per image with ConvNeXt backbone.
6.3 Ablation
# Prototypes CUB Accuracy
CBM 71.6
1 70.5
5 68.9
10 68.0
20 67.3
Table 2: Test Top-1 Accuracy on CUB with ResNet50 backbone with varying numbers of prototypes per concept.
We conduct an ablation on the number of prototypes per concept. Table 2 presents the test accuracies on CUB for HyperPgNet with ResNet50 backbone. The difference in task performance over the different model configurations is negligible. However, looking at some qualitative examples as depicted in Fig. 8 a difference in prototype quality can be observed. With fewer prototypes per concept, the model is more likely to also react to parts of the mage background in order to still achieve a low training error. With too many prototypes per concept, the latent space becomes too segmented. The prototypes start to focus on smaller subparts of the image subject, potentially decreasing the interpretability. The chosen configuration of 10 prototypes per concept seems to be a sweet spot regarding predictive performance and prototype quality for the experiments.
Figure 8: Qualitative Results for HyperPgNet with ResNet50 backbone and varying numbers of prototypes per concept.

### v3

@@ 3.2.1 Gaussian Prototypes
$=\frac{1}{(2\pi)^{\frac{D}{2}}|\bm{\mathrm{\Sigma}}|^{\frac{1}{2}}}$ $\exp\left(-\frac{1}{2}({\bm{z}}-\bm{{\mu}})^{\top}\bm{\mathrm{\Sigma}}^{-1}({\bm{z}}-\bm{{\mu}})\right).$ (4)
It is worth noting that the term inside the exponent is based on the $L_{2}$ distance between the latent vector ${\bm{z}}$ and the mean $\bm{{\mu}}$. This formulation as a PDF has several advantages: A) The similarity can be interpreted as the likelihood of being sampled from the Gaussian prototypes, which is more meaningful than a distance metric in a high-dimensional latent space. B) Prototypes can adapt their shape using a full covariance matrix, allowing different variances along various feature dimensions, offering more flexibility in shaping the latent space. However, learning a full covariance matrix increases computational requirements, especially with EM clustering.
(a) $\mu=1$
(b) $\mu=0.5$
(c) $\mu=0$
Figure 3: HyperPG prototypes learn a distribution of cosine similarities. They are parameterized by a learnable anchor vector $\bm{{\alpha}}$, scalar mean $\mu$ and scalar variance $\sigma^{2}$. HyperPG projects a Gaussian distribution of cosine similarities on the surface of a hypersphere, resulting in ring shaped activation patterns around the anchor vector.
Figure 4: Prototype Learning Architecture. The HyperPG module can be easily exchanged to other prototype formulations such as Euclidian or Cosine prototypes. HyperPG uses a Gaussian distribution as Density Estimator, but other PDFs are possible.
3.2.2 Gaussian Prototypes on the Hypersphere - HyperPG
Just as Gaussian prototypes provide a probabilistic formulation for Euclidean prototypes, we aim to develop a probabilistic formulation for hyperspherical prototypes based on the cosine similarity. We name this new formulation Prototypical Gaussians on the Hypersphere (HyperPG). HyperPG prototypes are defined as ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ with a directional anchor vector $\bm{{\alpha}}$, scalar mean similarity $\mu$ and scalar standard deviation (std) $\sigma$. HyperPG prototypes learn a $1D$ Gaussian distribution over the cosine similarities to the anchor vector $\bm{{\alpha}}$. Because the cosine similarity is bounded to $[-1,1]$, HyperPG’s similarity measure is defined as the PDF of the truncated Gaussian distribution within these bounds. Let $\mathcal{G}(x,\mu,\sigma)$ be the cumulative Gaussian distribution function. Then, HyperPG’s similarity measure based on the truncated Gaussian distribution is defined as
@@ 3.2.2 Gaussian Prototypes on the Hypersphere - HyperPG
Figure 3 illustrates the activations of HyperPG’s similarity function on the surface of a $3D$ hypersphere with anchor $\bm{{\alpha}}=(0,0,1)$, fixed std $\sigma=0.1$ and various mean values $\mu\in[0,1]$. The anchor $\bm{{\alpha}}$ defines a prototypical direction vector in latent space $\mathcal{Z}$, similar to other hyperspherical prototypes, and is visualized as a red arrow. The learned Gaussian distribution of cosine similarities is projected onto the hypersphere’s surface with std $\sigma$ governing the spread of the distribution, and mean $\mu$ the expected distance to the anchor $\bm{{\alpha}}$. For $\mu=1$, the distribution centers around the anchor, as the cosine similarity is $1$ if two vectors point in the same direction. This corresponds to the von Mises-Fisher (vMF) distribution (Appendix B). For $\mu=-1$, the distribution is the on opposite side of the hypersphere, as the cosine similarity is $-1$ for vectors pointing in opposite directions.
@@
For values of $11>\mu>-$ 1, the distribution forms a hollow ring around the anchor vector $\bm{{\alpha}}$. This occurs because the cosine similarity for these $\mu$ values expects the activating vectors to point in a different direction than the anchor, without specifying the direction. Imagining the interpolation between $\mu=1$ and $\mu=-1$, the probability mass moves from one pole of the hypersphere to the other, stretching like a rubber band over the surface. For $\mu=0$, the expected cosine similarity indicates that vectors with the highest activation are orthogonal to the anchor $\bm{{\alpha}}$. Since no specific direction is indicated, the entire hyperplanar segment orthogonal to the anchor has the highest activation.
These ring-like activation patterns would require an infinite mixture of prototype vectors pointing in all directions in this hyperplane. HyperPG achieves the same effect by learning only one prototype vector (the anchor) and just two additional scalar parameters. This significantly increases HyperPG’s representational power compared to Gaussian prototypes without increasing the computational complexity. This potential capability is also a major difference to the vMF distribution.
@@
HyperPG can be easily adapted to other probability distributions with additional desirable properties. Possible candidate distributions are elaborated on in Appendix B. Similarly, it is possible to exchange the cosine similarity to other similarity measures or functions, and learn an untruncated PDF over their output, making the HyperPG idea transferable to other manifolds and applications outside of prototype learning.
@@ 4 Training
The original ProtoPNet implementation uses three loss functions: a task specific loss like crossentropy for classification, a cluster loss to increase compactness within a class’s cluster, and a separation loss to increase distances between different prototype clusters.
@@ 4.1 Prototype Losses
where $s(\cdot,\cdot)$ is some similarity measure. The $L_{\mathrm{Clst}}$-Loss function increases compactness by increasing the similarity between prototypes ${\bm{p}}_{c}$ and latent embeddings ${\bm{z}}_{c}$ of class $c$ over all samples.
@@
The $L_{\mathrm{Sep}}$ function punishes high similarity values between a latent vector ${\bm{z}}_{c}$ of class c and prototypes ${\bm{p}}_{\neg c}$ not belonging to $c$, thereby separating the clusters in latent space. Please note, the original ProtoPNet paper uses a slightly different notation by working with the $L_{2}$ distance, instead of a similarity measure.
@@ 4.2 Multi-Objective Loss Function
$L=L_{\mathrm{CE}}+\lambda_{\mathrm{Clst}}L_{\mathrm{Clst}}+\lambda_{\mathrm{Sep}}L_{\mathrm{Sep}},$
where $L_{\mathrm{CE}}$ is the cross-entropy loss over network predictions and ground truth image labels. Our experiments uses $\lambda_{\mathrm{Clst}}=0.8$ and $\lambda_{\mathrm{Sep}}=0.08$ as proposed by ProtoPNet (Chen et al., 2019).
@@ 4.3 Network Architecture
For our experiments, we use the network architecture proposed by ProtoPNet (Chen et al., 2019), which we deem the basic deep prototype learning architecture for image classification. Figure 4 illustrates this network architecture with the HyperPG prototype module. In the encoding stage, a pretrained image encoder such as a ResNet model (He et al., 2016) serves as model backbone. Then, a convolution neck projects the high dimensional feature map of the backbone into the lower dimensional prototype space. This neck consists of two $1\times 1$ convolution layers with ReLU activation in between followed by a Sigmoid activation.
In the prototype stage, the prototype module computes the similarity scores. The original ProtoPNet uses Euclidian prototypes, which compute the $L_{2}$ distance between the encoded feature maps and the prototype parameters. Figure 4 shows the make up of the HyperPG module: First, the prototype module computes the cosine similarity of the learnable HyperPG anchors $\bm{{\alpha}}$ to the latent vectors produced by the neck. Second, a Density Estimation layer with learnable parameters mean $\mu$ and std $\sigma$ computes the Gaussian PDF over the activations of the previous layer. Future extensions could adapt both components independently, for example by implementing a hyperbolic similarity measure or a multi-modal probability distribution.
The final classification stage produces the output logits based on the prototype module’s similarity scores. ProtoPNet uses a single linear layer due to its inherent interpretability.
4.3.1 Comparison to Recent Works
Our experiments use only the basic ProtoPNet architecture and focus on the effect of different prototype formulations. This section provides a short overview of the required changes to this basic architecture in order to implement the models described in Table 1. LucidPPN (Pach et al., 2024) replaces the encoding stage with a hybrid encoder specializing in color-only and texture-only feature extraction. Deformable ProtoPNet (Donnelly et al., 2021) makes changes towards the input of the prototype module to allow for dynamic spatial arrangements of the prototypes, i.e., “deformations”. ProtoPool (Rymarczyk et al., 2022) makes changes towards the output of the prototype module, by also learning how to dynamically pool the different prototype activations. ProtoTree (Nauta et al., 2021) and ProtoKNN (Ukai et al., 2023) propose to use a different inherently interpretable model for the classification stage. ProtoPFormer (Xue et al., 2024) builds exclusively on a Vision Transformer backbone and uses different network branches for the global information encoded in the class token and the local information in the image tokens.
@@ 5 Experiments
The experiments were performed on CUB-200-2011 (CUB) with 200 bird species (Wah et al., 2011) with ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) backbones. Later experiments also explore the Stanford Cars (Krause et al., 2013) and Oxford Flowers (Nilsback & Zisserman, 2008) datasets. Implementation details such as data preprocessing or model implementation are provided in Appendix D.
Table 2: Mean Top-1 Test Accuracies and Std for different prototype formulations with ResNet50 Backbones averaged over 3 random seeds.
ResNet50 Dense121 ViT-B16
Baseline 79.2 $\pm$ 0.2 75.5 $\pm$ 0.2 78.5 $\pm$ 0.8
Euclidian (ProtoPNet) 35.3 $\pm$ 1.4 33.6 $\pm$ 3.0 9.4 $\pm$ 0.4
Gaussian 61.1 $\pm$ 0.2 45.9 $\pm$ 5.1 4.2 $\pm$ 1.3
Cosine 71.7 $\pm$ 0.4 68.7 $\pm$ 0.5 69.2 $\pm$ 3.9
HyperPG (Ours) 74.3 $\pm$ 0.6 70.7 $\pm$ 0.3 72.9 $\pm$ 0.1
@@ 5.1 Quantitative Results
Table 2 reports the mean top-1 accuracy and standard deviation for the tested models on the CUB dataset with backbone models ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) pretrained on ImageNet. Please note that in contrast to prior work, we do not report scores from model ensembles or using pretrained iNaturalist weights. The black-box baselines perform the best across all experiments, which we see as a theoretical upper bound for the performance. The classical Euclidean prototypes perform the worst across all experiments. A first model improvement can be achieved by switching to probabilistic Gaussian prototypes. The hyperspherical Cosine prototypes further improve the performance. HyperPG combines the advantages of Gaussian and Cosine prototypes, further closing the gap to the upper bound presented by the black-box baselines. Notably, both Euclidian and Gaussian prototypes are unable to learn on the latent features of ViT-B16. We theorize that the attention mechanism of the Transformer architecture produces a near-hyperspherical latent space. This highlights the importance of choosing the right prototype formulation for the used backbone. For example, language aligned models such as (Radford et al., 2021) use the cosine similarity for their alignment. We expect that only hyperspherical prototypes can produce meaningful results in this setting.
Figure 5: CUB-200-2011 Test Accuracy per Epoch with ResNet50 backbone. Mean and std over 3 random seeds.
Figure 5 shows the test accuracy over epochs for CUB with ResNet50 backbone. The standard Euclidean prototypes take the longest time (200 Epochs) until convergence while also achieving the lowest accuracy. By changing only the prototype formulation to a Gaussian one, the model achieves higher performance and converges at 100 epochs. Both the cosine and HyperPG formulation outperform the other formulations in model accuracy and training time, requiring only 30 training epochs.
5.2 Ablation: Number of Prototypes
Figure 6: Ablation: Test Accuracy on CUB with different numbers of prototypes per class.
We conduct an ablation study over the number of prototypes per class and present the test accuracy scores in Figure 6. Notably, the predictive performance of Euclidean and Gaussian prototypes degrades when more prototypes are added. The prototype formulations with a hyperspherical structure, Cosine and HyperPG prototypes converge to similar performance levels regardless of the number of prototypes per class.
5.3 Ablation: Number of Dimensions
We conduct an ablation study over the number of prototype dimensions and present the test accuracy scores in Figure 7. Hyperspherical prototype formulations such as Cosine and HyperPG are robust regarding the choice of prototype dimensions, although HyperPG performs slightly better with increased dimensionality. The Gaussian prototypes, like HyperPG prototypes, can adapt their variance $\sigma^{2}$ to deal with the increase in dimensionality. The performance of Euclidean prototypes suffers most with changes in dimensionality.
Figure 7: Ablation: Test Accuracy on CUB with different numbers of prototype dimensions.
5.4 Ablation: Additional Datasets
We perform additional experiments on two different datasets. Stanford Cars (Cars) with 196 car models (Krause et al., 2013) and Oxford Flowers (Flowers) with 102 different species (Nilsback & Zisserman, 2008). The relative differences in performance seem to hold across the different datasets, with Euclidean prototypes performing the worst. Interestingly, Gaussian prototypes are able to match the performance of the hyperspherical formulations more closely.
Table 3: Top-1 Test Accuracy with Resnet50 Backbone.
Cars Flowers
Baseline 85.2 90.0
Euclidean (ProtoPNet) 36.2 15.0
Gaussian 73.0 74.8
Cosine 79.6 85.5
HyperPG (Ours) 79.0 87.8
6 Interpretability Analysis
We perform a qualitative analysis of the different prototype formulations with ResNet50 backbone (see Figure 8). One common visualization technique originally proposed by ProtoPNet is a a direct overlay of the prototype activation map (PAM). However, this visualization technique has several shortcomings, which we discuss in more detail in Appendix A. In order to offer an alternative to the PAM visualization, we also present GradCAM (Selvaraju et al., 2017) overlays.
Both Euclidian and HyperPG prototypes do not actually learn class-specific prototypes, in contrast to Gaussian or Cosine prototypes. Interestingly, all prototype formulations seem to activate in similar regions in the test image, but differ with the known training examples. HyperPG could already leverage class-shared prototypes, which would explain its robustness to changes in the number of prototypes per class.
There seems to be no qualitative difference between the visualizations produced with the various prototype formulations using the CNN based backbones. As expected, the ViT backbone (see Figure 9) suffers from the “distraction problem” as observed by ProtoPFormer (Xue et al., 2024). We present additional examples in Appendix A.
Refer to caption Refer to caption
(a) Euclidean (R50)
Refer to caption Refer to caption
(b) Gaussian (R50)
Refer to caption Refer to caption
(c) Cosine (R50)
Refer to caption Refer to caption
(d) HyperPG (R50)
Figure 8: Left: Test Image. Right: 3 closest Training Patches to the Prototype. Top Row: PAM visualization. Bottom Row: GradCAM visualization. R50: Resnet50 Backbone. ViT: ViT-B16 Backbone.
Figure 9: Cosine Protoypes with ViT. The overlayed heatmap show the distraction problem caused by the attention mechanism.